OpenStudy (anonymous):

5.7333333333 repeat as a fraction I.E 5 6/8

OpenStudy (anonymous):

@Hero

hero (hero):

Okay, basically for this first you'll need to do some side work.

OpenStudy (anonymous):

like what?

hero (hero):

Let 100x = 73.333333... 10x = 7.333333... Then subtract 100x - 10x = 73.333333.... - 7.333333... 90x = 66 x = 66/90 x = 11/15 But don't forget about the 5 5 + 11/15

hero (hero):

I don't know if that's how you were taught how to do this or not.

OpenStudy (anonymous):

5 11/12???

OpenStudy (anonymous):

its not a choice only 5 11/15

hero (hero):

Where'd you get 5 + 11/12 from?

OpenStudy (anonymous):

like u said

hero (hero):

Check with your calculator. You'll see for yourself.

OpenStudy (anonymous):

ok thanks!

hero (hero):

5 11/15 is the same as 5 + 11/15. Remember what I explained to you earlier?

hero (hero):

I'm curious to know how you were taught to do this. If there's another method, which I'm sure there is, I'd like to know.